# CRANE: Causal Relevance Analysis of Language-Specific Neurons in Multilingual Large Language Models

### v1

@@ Abstract
Multilingual large language models (LLMs) achieve strong performance across languages, yet how language capabilities are organized at the neuron level remains poorly understood. Prior work has identified language-related neurons mainly through activation-based heuristics, which conflate language preference with functional importance. Prior work has identified language-related neurons mainly through activation-based heuristics, which conflate language preference with functional importance. We propose CRANE, a relevance-based analysis framework that redefines language specificity in terms of functional necessity, identifying language-specific neurons through targeted neuron-level interventions. CRANE characterizes neuron specialization by their contribution to language-conditioned predictions rather than activation magnitude. Our implementation will be made publicly available. Neuron-level interventions reveal a consistent asymmetric pattern: masking neurons relevant to a target language selectively degrades performance on that language while preserving performance on other languages to a substantial extent, indicating language-selective but non-exclusive neuron specializations. Experiments on English, Chinese, and Vietnamese across multiple benchmarks, together with a dedicated relevance-based metric and base-to-chat model transfer analysis, show that CRANE isolates language-specific components more precisely than activation-based methods.
@@ 1 Introduction
A growing line of work has attempted to identify language-specific neurons, often based on activation statistics or language-conditioned probing. Earlier studies on recurrent neural networks examined gated units as mechanisms for information storage and control (Hochreiter and Schmidhuber, 1997; Elman, 1990), while later work analyzed intrinsic linguistic knowledge in sequential models (Lu et al., 2018). More recently, neuron-level analyses in LLMs have focused on activation distributions and activation probabilities (Tang et al., 2024), as well as finer-grained neuron identification approaches (Zhang et al., 2025; Sachan et al., 2025). These studies reveal statistical regularities associated with language, but largely assume that statistical correlation reflects functional importance.
@@ 2 Related Work
For multilingual models, prior work studies cross-lingual alignment and shared representations using probing and representation similarity analysis Karthikeyan et al. (2020); Blevins et al. (2022). More recent studies investigate language-specific or language-selective neurons by analyzing activation statistics or frequency, exemplified by LAPE Tang et al. (2024) and related neuron-level methods Zhang et al. (2025); Sachan et al. (2025). While these methods uncover language-related structure, they typically equate activation selectivity with functional importance. As noted in prior work, such selectivity does not guarantee that intervening on the identified neurons induces targeted language degradation Tang et al. (2024).
@@
A growing line of mechanistic interpretability emphasizes causal validation via intervention. Methods such as activation patching and causal tracing intervene on internal activations to localize computations responsible for model behavior Meng et al. (2022); Zhang et al. (2024). These approaches motivate defining interpretability in terms of functional necessity rather than correlational signals. Our work aligns with this perspective and applies intervention-based validation to study language specialization in multilingual LLMs.
@@ 4.3 Tasks and Evaluation
We evaluate language understanding using standard language-specific benchmarks: MMLU for English (Hendrycks et al., 2021), C-Eval for Chinese (Huang et al., 2023), and Belebele for Vietnamese (Bandarkar et al., 2023). All NLU evaluations are conducted using a unified evaluation framework, lm-evaluation-harness (Biderman et al., 2024), with consistent prompting and decoding settings across languages and methods.
@@
We evaluate open-ended generation using an English/Chinese/Vietnamese question set following the LAPE protocol. Model outputs are scored on a 1–10 scale by an LLM-as-a-judge based on GPT-4o(Hurst et al., 2024). The same evaluation protocol is applied before and after neuron-level intervention to ensure direct comparability.
@@ 4.4 Inference and Implementation Details
Inference and evaluation are performed using vLLM(Kwon et al., 2023). We use greedy decoding with repetition_penalty=$1.1$, max_new_tokens=20482048, and a maximum context length of $8192$. For each identified neuron set, we evaluate all tasks both before and after masking to ensure direct comparison of intervention effects.

### v2

@@ Abstract
Multilingual large language models (LLMs) achieve strong performance across languages, yet how language capabilities are organized at the neuron level remains poorly understood. Prior work has identified language-related neurons mainly through activation-based heuristics, which conflate language preference with functional importance. We propose CRANE, a relevance-based analysis framework that redefines language specificity in terms of functional necessity, identifying language-specific neurons through targeted neuron-level interventions. CRANE characterizes neuron specialization by their contribution to language-conditioned predictions rather than activation magnitude. Our implementation will be made publicly available. Neuron-level interventions reveal a consistent asymmetric pattern: masking neurons relevant to a target language selectively degrades performance on that language while preserving performance on other languages to a substantial extent, indicating language-selective but non-exclusive neuron specializations. Experiments on English, Chinese, and Vietnamese across multiple benchmarks, together with a dedicated relevance-based metric and base-to-chat model transfer analysis, show that CRANE isolates language-specific components more precisely than activation-based methods.
@@ 1 Introduction
A growing line of work has attempted to identify language-specific neurons, often based on activation statistics or language-conditioned probing. Earlier studies on recurrent neural networks examined gated units as mechanisms for information storage and control (Hochreiter and Schmidhuber, 1997; Elman, 1990), while later work analyzed intrinsic linguistic knowledge in sequential models (Qian et al., 2016). More recently, neuron-level analyses in LLMs have focused on activation distributions and activation probabilities (Tang et al., 2024), as well as finer-grained neuron identification approaches (Dai et al., 2022; Zhang et al., 2025). These studies reveal statistical regularities associated with language, but largely assume that statistical correlation reflects functional importance.
@@ 2 Related Work
For multilingual models, prior work studies cross-lingual alignment and shared representations using probing and representation similarity analysis Karthikeyan et al. (2020); Blevins and others (2022). More recent studies investigate language-specific or language-selective neurons by analyzing activation statistics or frequency, exemplified by LAPE Tang et al. (2024) and related neuron-level methods Zhang et al. (2025). While these methods uncover language-related structure, they typically equate activation selectivity with functional importance. As noted in prior work, such selectivity does not guarantee that intervening on the identified neurons induces targeted language degradation Tang et al. (2024).
@@
A growing line of mechanistic interpretability emphasizes causal validation via intervention. Methods such as activation patching and causal tracing intervene on internal activations to localize computations responsible for model behavior Meng et al. (2022); Zhang and others (2024). These approaches motivate defining interpretability in terms of functional necessity rather than correlational signals. Our work aligns with this perspective and applies intervention-based validation to study language specialization in multilingual LLMs.
@@ 4.3 Tasks and Evaluation
We evaluate language understanding using standard language-specific benchmarks: MMLU for English (Hendrycks and others, 2021), C-Eval for Chinese (Huang et al., 2023), and Belebele for Vietnamese (Bandarkar et al., 2023). All NLU evaluations are conducted using a unified evaluation framework, lm-evaluation-harness (Biderman et al., 2024), with consistent prompting and decoding settings across languages and methods.
@@
We evaluate open-ended generation using an English/Chinese/Vietnamese question set following the LAPE protocol. Model outputs are scored on a 1–10 scale by an LLM-as-a-judge based on GPT-4o(Hurst and others, 2024). The same evaluation protocol is applied before and after neuron-level intervention to ensure direct comparability.
@@ 4.4 Inference and Implementation Details
Inference and evaluation are performed using vLLM(Kwon and others, 2023). We use greedy decoding with repetition_penalty=$1.1$, max_new_tokens=$2048$, and a maximum context length of $8192$. For each identified neuron set, we evaluate all tasks both before and after masking to ensure direct comparison of intervention effects.